# Çukurova Üniversitesi Rektörlüğü 01250 Sarıçam, Adana, Turkey

### v1

@@ 5 Performance of DRN with beam test data
We tested the DRNs with real data collected with a negatively charged pion beam at the CERN SPS in 2018. The apparatus is described in Section 2. Event samples of pions were collected with beam energy settings of 20, 50, 80, 100, 120, 200, 250 and 300 GeV. The expected purity of the beam was 80–90% with the balance consisting mostly of muons, based on a detailed simulation of the beam line ShubhamThesis . The sample size for each energy setting ranged from 80 k--100 k events collected and 40 k--80 k events are available after fiducial and other quality selections for further analysis. The selection criteria are described in Ref. bib.hgcal-2018-pions . Approximately 100k events were simulated with Geant4 for each of the eight energy points using the same geometry and the same physics lists that were used for the training samples. The same event selection criteria applied to events in the data were applied to the simulated events. These fixed energy samples are not used for model training.
For comparison of the detector response between beam test data and simulation, we used the DRN (E,x,y,z) model which had given the best results. We also used WS method for comparison. It is identical to the one described in the Ref. bib.hgcal-2018-pions and is re-evaluated for the configuration with reduced number AHCAL layers that we used. Comparisons of the energy distribution found using the DRN method in the data and in the simulation are shown in Fig. 12 for beam energies of 20, 50, 100 and 200 GeV, with the model accurately reproducing the overall shape of the measured energy distributions. We plot in Fig. 13 the energy resolution and the energy response as a function of incident energy found with the DRN(E,x,y,z) model and with the WS method.
@@
To understand why the DRN significantly improves over the WS method, we note that due to the high granularity of the detector the DRN is able to account for the event-by-event variations in energy sharing between the electromagnetic and hadronic components. This is illustrated by distribution of energy reconstructed in simulated events by DRN (E,x,y,z) model and the WS method as a function of the fraction of incident energy carried by the $\pi^{0}$s for 50 GeV and 200 GeV pions as shown in Fig. 14. The fraction of energy carried by $\pi^{0}$s is obtained directly from Geant4.
Figure 14: Energy reconstructed using WS method (top panel) and DRN(E,x,y,z) model (lower panel) for pions of energy 50 GeV (left) and 200 GeV (right) as a function of fraction of energy carried by $\pi^{0}$s in a given shower.
@@ 6 Expected performance using the full HGCAL geometry
Due to its limited size, energy losses from the prototype calorimeter were significant. This will not be the case with the final HGCAL that will be installed in the CMS experiment. To quantify the effect of potential leakage of energy and to validate the performance of the GNN-based algorithm in a realistic setup, we tested the model using a sample of pions generated with a standalone full-geometry setup simulated using Geant4. The simulation did not include the upstream material that will be present in the final detector. It does, however, include a detector configuration of sensor modules with areas 0.5 $cm^{2}$ and 1 $cm^{2}$ and low density modules as shown in Fig. 15, which reflects better the final HGCAL configuration than the one described in Ref. bib.cms-hgcalphase2-tdr . In this figure, $Z$ is the horizontal distance from the interaction point and $\rho$ is the radial distance from the beamline.
@@
We simulated the pion showers entering at $\eta=2$ in this setup corresponding to $\rho$=800 mm in Fig. 16 and separately trained the model with the corresponding inputs. As in the simulation of the beam test detector setup, the signal deposited in each detector channel is converted to the unit of number of MIPs by factors determined with simulated 200 GeV muons with the digitization and electronics noise superimposed on the signal. As before, rechits with energy deposits of less than 0.5 MIP are rejected. Three million charged pion events were simulated with the energies distributed uniformly from 10 to 350 GeV. We use the rechits energies and the spatial coordinates to train the models DRN (E), DRN (E,z) and DRN (E,x,y,z). The resolutions and responses obtained with these three models are shown in Fig. 16. We observed that adding information about the transverse coordinates of rechits has little impact on the predictions based on the DRN which is trained on input features comprising rechit energies and depth. This result can be understood as being due to the full shower containment in this extended geometry, unlike the smaller prototype where there was energy leakage. The model incorporates additional information from the transverse distribution of rechits provided as $x$ and $y$ positions. For a fully contained shower, the model learns the true energy based on the highly granular distribution of energies of rechits and their depth in the detector.
@@
Figure 17: Distributions of of energy reconstructed in full HGCAL and beam test setups using DRN(E,x,y,z) for pions with energies 20 GeV (left) and 200 GeV (right).
@@ 7 Summary
Acknowledgements.
We thank the technical and administrative staffs at CERN and at other CMS institutes for their contributions to the success of the CMS effort. We acknowledge the enduring support provided by the following funding agencies and laboratories: BMBWF and FWF (Austria); CERN; CAS, MoST, and NSFC (China); MSES and CSF (Croatia); CEA, CNRS/IN2P3 and P2IO LabEx (ANR-10-LABX-0038) (France); SRNSF (Georgia); BMBF, DFG, and HGF (Germany); GSRT (Greece); DAE and DST (India); MES (Latvia); MOE and UM (Malaysia); MOS (Montenegro); PAEC (Pakistan); FCT (Portugal); NSTC (Taipei); ThEP Center, IPST, STAR, and NSTDA (Thailand); TUBITAK and TENMAK (Turkey); STFC (United Kingdom); and DOE (USA). The authors acknowledge the Minnesota Supercomputing Institute (MSI) at the University of Minnesota and National Supercomputing Mission (NSM) at the Indian Institute of Science Education and Research, Pune for providing resources that contributed to the research results reported within this paper.

### v3

@@ 5 Performance of DRN with beam test data
We tested the DRNs with real data collected with a negatively charged pion beam at the CERN SPS in 2018. The apparatus is described in Section 2. Event samples of pions were collected with beam energy settings of 20, 50, 80, 100, 120, 200, 250 and 300 GeV. The expected purity of the beam was 80–90% with the balance consisting mostly of muons, based on a detailed simulation of the beam line [29]. The sample size for each energy setting ranged from 80 k--100 k events collected and 40 k--80 k events are available after fiducial and other quality selections for further analysis. The selection criteria are described in Ref. [11]. Approximately 100k events were simulated with Geant4 for each of the eight energy points using the same geometry and the same physics lists that were used for the training samples. The same event selection criteria applied to events in the data were applied to the simulated events. These fixed energy samples are not used for model training.
For comparison of the detector response between beam test data and simulation, we used the DRN (E,x,y,z) model which gave the best results. We also used the WS method for comparison. It is identical to the one described in the Ref. [11] and is re-evaluated for the configuration with a reduced number AHCAL layers that we used. Comparisons of the energy distribution found using the DRN method in the data and in the simulation are shown in Fig. 12 for beam energies of 20, 50, 100 and 200 GeV, with the model accurately reproducing the overall shape of the measured energy distributions. We plot in Fig. 13 the energy resolution and the energy response as a function of incident energy found with the DRN(E,x,y,z) model and with the WS method.
@@
To understand why the DRN significantly improves over the WS method, we note that due to the high granularity of the detector the DRN is able to account for the event-by-event variations in energy sharing between the electromagnetic and hadronic components. This is illustrated by the distribution of energy reconstructed in simulated events by the DRN (E,x,y,z) model and the WS method as a function of the fraction of incident energy carried by the $\pi^{0}$s for 50 GeV and 200 GeV pions as shown in Fig. 14. The fraction of energy carried by $\pi^{0}$s is obtained directly from Geant4.
Figure 14: Energy reconstructed using WS method (top panel) and DRN(E,x,y,z) model (lower panel) for pions of energy 50 GeV (left) and 200 GeV (right) as a function of the fraction of energy carried by $\pi^{0}$s in a given shower.
@@ 6 Expected performance using the full HGCAL geometry
Due to its limited size, energy losses from the prototype calorimeter were significant. This will not be the case with the final HGCAL that will be installed in the CMS experiment. To quantify the effect of potential leakage of energy and to validate the performance of the GNN-based algorithm in a realistic setup, we tested the model using a sample of pions generated with a standalone full-geometry setup simulated using Geant4. The simulation did not include the upstream material that will be present in the final detector. It does, however, include a detector configuration of sensor modules with areas 0.5 $cm^{2}$ and 1 $cm^{2}$ and low density modules as shown in Fig. 15, which reflects the final HGCAL configuration better than the one described in Ref. [2]. In this figure, $Z$ is the horizontal distance from the interaction point and $\rho$ is the radial distance from the beamline.
@@
We simulated the pion showers entering at $\eta=2$ in this setup corresponding to $\rho$=800 mm in Fig. 16 and separately trained the model with the corresponding inputs. As in the simulation of the beam test detector setup, the signal deposited in each detector channel is converted to the unit of number of MIPs by factors determined with simulated 200 GeV muons with the digitization and electronics noise superimposed on the signal. As before, rechits with energy deposits of less than 0.5 MIP are rejected. Three million charged pion events were simulated with the energies distributed uniformly from 10 to 350 GeV. We use the rechits energies and the spatial coordinates to train the models DRN (E), DRN (E,z), and DRN (E,x,y,z). The resolutions and responses obtained with these three models are shown in Fig. 16. We observed that adding information about the transverse coordinates of rechits has little impact on the predictions based on the DRN which is trained on input features comprising rechit energies and depth. This result can be understood as being due to the full shower containment in this extended geometry, unlike the smaller prototype where there was energy leakage. The model incorporates additional information from the transverse distribution of rechits provided as $x$ and $y$ positions. For a fully contained shower, the model learns the true energy based on the highly granular distribution of energies of rechits and their depth in the detector.
@@
Figure 17: Distributions of energy reconstructed in full HGCAL and beam test setups using DRN(E,x,y,z) for pions with energies 20 GeV (left) and 200 GeV (right).
@@ 7 Summary
Acknowledgments
We thank the technical and administrative staffs at CERN and at other CMS institutes for their contributions to the success of the CMS effort. We acknowledge the enduring support provided by the following funding agencies and laboratories: BMBWF and FWF (Austria); CERN; CAS, MoST, and NSFC (China); MSES and CSF (Croatia); CEA, CNRS/IN2P3 and P2IO LabEx (ANR-10-LABX-0038) (France); SRNSF and GTU (Georgia); BMBF, DFG, and HGF (Germany); GSRT (Greece); DAE and DST (India); MES (Latvia); MOE and UM(Malaysia); MOS (Montenegro); PAEC (Pakistan); FCT (Portugal); NSTC (Taipei); ThEP Center, IPST, STAR, and NSTDA (Thailand); TUBITAK and TENMAK (Turkey); STFC (United Kingdom); and DOE (USA). The authors acknowledge the Minnesota Supercomputing Institute (MSI) at the University of Minnesota and the National Supercomputing Mission (NSM) at the Indian Institute of Science Education and Research, Pune for providing resources that contributed to the research results reported within this paper.